# What Sensors See, What People Feel: Exploring Subjective Collaboration Perception in Mixed Reality

### v2

@@ 4.3 Performance via Sensor-based Indicators
| Sensor Metric | Target Normality (Indiv. / Group) | Metric Normality (Indiv. / Group) | Correlation (Indiv. / Group) | p-Value (Indiv. / Group) | Test Used (Indiv. / Group) |
| Shared Gaze Overlap Frequency | Spearman / Spearman | ||||
| Proximity Duration | (R2) | Linear / Linear | |||
| Image Grabs Variance | Spearman / Spearman | ||||
| Label Overrides Variance | Spearman / Pearson | ||||
@@ 4.4 Experience via Sensor-based Indicators
| Sensor Metric | Target Normality (Indiv. / Group) | Metric Normality (Indiv. / Group) | Correlation) (Indiv. / Group) | p-Value (Indiv. / Group) | Test Used (Indiv. / Group) |
| Target: Presence (PQ + IPQ) | |||||
| Shared Attention (%) | 0.8677 / 0.7494 | 0.3722 / 0.9363 | 0.0417 / 0.4802 | 0.7783 / 0.1141 | Pearson / Pearson |
| Jointly Viewed Images | 0.8677 / 0.7494 | 0.2048 / 0.3740 | -0.0820 / -0.4956 | 0.5795 / 0.1013 | Pearson / Pearson |
| Mean Pairwise Distance | 0.8677 / 0.7494 | 0.5384 / 0.7793 | 0.0736 / -0.0282 | 0.6190 / 0.9307 | Pearson / Pearson |
| Proximity Duration | 0.8677 / 0.7494 | 7.3267e-07 / 0.0650 | -0.0536 / 0.2626 | 0.7173 / 0.4097 | Spearman / Pearson |
| Variance in Image Grabs | 0.8677 / 0.7494 | 7.9738e-11 / 7.2687e-05 | 0.0083 / -0.1049 | 0.9555 / 0.7456 | Spearman / Spearman |
| Target: Cognitive Load (NASA TLX) | |||||
| Shared Gaze Overlap | 0.0225 / 0.3828 | 7.3990e-08 / 0.0014 | 0.3223 / 0.2872 | 0.0255 / 0.3654 | Spearman / Spearman |
| Pairwise Distance | 0.0225 / 0.3828 | 0.5384 / 0.7793 | -0.1777 / -0.2020 | 0.2268 / 0.5289 | Spearman / Pearson |
| Variance in Image Grabs | 0.0225 / 0.3828 | 7.9738e-11 / 7.2687e-05 | 0.1261 / 0.2522 | 0.3933 / 0.4291 | Spearman / Spearman |
@@ 5 Discussion
Sensory behavioral features were weakly associated with subjective states such as the presence or cognitive load. A near-significant individual-level link was observed between cognitive load and shared gaze frequency, but it lacked strength and consistency at the group level. Presence did not correspond clearly with proximity or gaze patterns, suggesting a lack of visible manifestation of internal states in this context. Likewise, no significant relationship emerged between cognitive load and participation balance. H3 is partially accepted due to some individual-level signal, but not consistently. We reject H3.1, H3.2, and H3.3 due to a lack of strong evidence connecting presence or cognitive load to group-level behaviors. Although disappointing at first glance, this result replicates concerns by others [9, 25] that objective signals often fail to capture subjective experience unless paired with richer data sources.
@@ 5.1 Implications
Our results show that while sensed behavior reliably maps onto collaborative perception and task efficiency, it falls short of capturing deeper subjective experiences, such as presence or cognitive availability, in its current form. This reinforces that not all internal states are externally observable (3), even in high-fidelity sensing environments. This absence of mapping is not a flaw it is a vital empirical finding that delineates what current MR sensing can (and importantly cannot) reveal about subjective experience. While prior works have treated gaze or physiological synchrony as proxies for presence or co-presence [28], our results underscore their limitations without deeper context. Our findings align with recent studies showing that even rich physiological or behavioral data often fail to reliably predict internal states unless complemented with richer modalities, such as neural, situational, or self-report inputs [40, 25, 29]. To meaningfully comprehend experience from sensor data, some intermediary techniques are needed to connect high-level behavioral indicators with low-level experiential states. Designers should treat behavioral signals as interpretive anchors, not definitive measures of internal states, reinforcing the cautious approach seen in immersive cognitive sensing research, and instead consider combining sensing with self-reports or adaptive prompting. Hybrid models that integrate system-logged data with lightweight subjective inputs could enable more robust personalization and responsiveness in collaborative MR.
@@ 6 Limitations and Future Work
This study offers early insight into the relationship between sensed group behavior, subjective perception, and task performance in collaborative MR settings. At the same time, several limitations should be considered, some by design, others presenting opportunities for future work.
First, we employed a singular task in one experimental condition to distinctly observe collaboration dynamics in a regulated environment. The open-ended image sorting task was chosen due to its flexibility and inherent coordination requirements, which are ideal for observing genuine group behavior. Despite this, the controlled environment might not reflect the complete range of variability in time-sensitive or specialized tasks. Future research should investigate various task types and adaptive system conditions to generalize beyond this setup. The number of groups restricts the statistical power of analyses conducted at the group level. Despite a relatively high participant count, the nature of small-group interactions limits group-level observations. Nonetheless, the sample size was adequate to reveal consistent and interpretable patterns across both individual and group analyses. Future research could build upon this by employing larger-scale deployments or multi-session studies.
@@
Group composition was not experimentally manipulated; participants had the autonomy to self-select or be randomly assigned to groups. This decision sought to observe genuine interaction dynamics instead of creating artificial teams. Nonetheless, pre-existing familiarity among participants might impact conversation, shared focus, and comfort. Although not measured in this study, future research could quantify prior familiarity to better understand its effects.
@@
In addition, we focus on interaction signals such as gaze, proximity, and voice activity, which could be tracked using standard MR hardware. This choice aligned with our aim to create practical, lightweight sensing systems. Yet, using only headset sensors restricted us from accessing detailed affective, physiological, or facial data. Future research should investigate multimodal enhancements for richer sensing, such as emotion recognition or biometric monitoring. Finally, while our framework bridges sensor data and subjective reflection, it cannot fully capture internal states such as frustration, motivation, or attention lapses. These nuanced experiences often remain invisible to behavioral sensing alone. Our study highlights the gap between observable behavior and internal experience, reinforcing the need for interpretive models that integrate multiple modalities and participant feedback.
In sum, the study was purposefully scoped to explore collaborative sensing in MR using a task and instrumentation that balance ecological validity and control. The observed patterns suggest promising directions for expanding system awareness and adaptability while motivating future research to test across conditions, sensor modalities, and longer-term collaboration scenarios.
@@ 7 Conclusion
This study examines the alignment between objective sensor data and subjective perception of experiences in MR collaboration. The confirmation of H1 and H2 indicates that sensor-derived metrics, such as shared gaze frequency, reliably reflect perceived collaboration and task performance. Nonetheless, the partial validation of H3 reveals limitations: complex subjective states, such as presence and cognitive load, show limited correlation with objective measures of behavioral data, suggesting that internal experiences are difficult to capture solely through external observations. Future investigations should focus on the interaction between objective and subjective metrics across varied conditions, sensor types, and collaboration contexts.
We situate this work in the growing body of research studying design-responsive, human-aware MR systems. By grounding system- and sensor-level observations in subjective experience, we move toward sensing models that prioritize users’ perspectives. This paper contributes a conceptual model, a validated experimental setup, and empirical findings illuminating where sensor data and user experience align and where they diverge. Through this lens, we take a step toward MR environments that do not just support collaboration but sense it in the ways users feel it.
@@ Acknowledgements.
- [1] Hololens 2. https://www.microsoft.com/en-us/hololens/hardware. Accessed Sep-2024.
- [2] Apple Vision Pro. {https://www.apple.com/apple-vision-pro/}, 2023.
@@
- [4] H. Bai, P. Sasikumar, J. Yang, and M. Billinghurst. A user study on mixed reality remote collaboration with eye gaze and hand gesture sharing. In Proceedings of the 2020 CHI conference on human factors in computing systems, pp. 1–13, 2020.
- [5] J. N. Bailenson, A. C. Beall, and J. Blascovich. Gaze and task performance in shared virtual environments. The Journal of Visualization and Computer Animation, 13(5):313–320, 2002. _eprint: https://onlinelibrary.wiley.com/doi/pdf/10.1002/vis.297. doi: 10 . 1002/vis . 297
- [6] J. N. Bailenson, A. C. Beall, J. Loomis, J. Blascovich, and M. Turk. Transformed social interaction: Decoupling representation from behavior and form in collaborative virtual environments. Presence: Teleoperators & Virtual Environments, 13(4):428–441, 2004.
- [7] J. N. Bailenson, N. Yee, J. Blascovich, A. C. Beall, N. Lundblad, and M. Jin. The use of immersive virtual reality in the learning sciences: Digital transformations of teachers, students, and social context. The journal of the learning sciences, 17(1):102–141, 2008.
- [8] R. Bakeman and L. B. Adamson. Coordinating Attention to People and Objects in Mother-Infant and Peer-Infant Interaction. Wiley, Society for Research in Child Development, 55(4):1278–1289, 1984.
- [9] A. Bayro, Y. Ghasemi, and H. Jeong. Subjective and objective analyses of collaboration and co-presence in a virtual reality remote environment. In 2022 IEEE Conference on Virtual Reality and 3D User Interfaces Abstracts and Workshops (VRW). IEEE, 2022.
@@
- [11] D. Bohus and E. Horvitz. Facilitating multiparty dialog with gaze, gesture, and speech. In International Conference on Multimodal Interfaces and the Workshop on Machine Learning for Multimodal Interaction, ICMI-MLMI ’10. Association for Computing Machinery, New York, NY, USA, 2010. doi: 10 . 1145/1891903 . 1891910
@@
- [40] N. Numan and A. Steed. Exploring User Behaviour in Asymmetric Collaborative Mixed Reality. In Proceedings of the 28th ACM Symposium on Virtual Reality Software and Technology, VRST ’22. New York, NY, USA, 2022.
@@
- [47] Z. Qu, R. Byrne, and M. Gorlatova. “looking” into attention patterns in extended reality: An eye tracking-based study. In 2024 IEEE International Symposium on Mixed and Augmented Reality (ISMAR), 2024.
@@
- [49] J. A. Russell. A circumplex model of affect. Journal of Personality and Social Psychology, 39(6):1161–1178, Dec. 1980. doi: 10 . 1037/h0077714
@@
- [66] G. Vissers and B. Dankbaar. Knowledge and Proximity. European Planning Studies, 21(5):700–721, May 2013. doi: 10 . 1080/09654313 . 2013 . 734459
@@
- [70] Y. Yang, T. Dwyer, M. Wybrow, B. Lee, M. Cordeil, M. Billinghurst, and B. H. Thomas. Towards Immersive Collaborative Sensemaking. Proceedings of the ACM on Human-Computer Interaction, 6:722–746, Nov. 2022.

### v3

@@ 4.3 Performance via Sensor-based Indicators
| Sensor Metric | Target Normality (Indiv. / Group) | Metric Normality (Indiv. / Group) | Correlation (Indiv. / Group) | p-Value (Indiv. / Group) | q-Value (FDR) (Indiv. / Group) | Test Used (Indiv. / Group) |
| Shared Gaze Overlap Frequency | Spearman / Spearman | |||||
| Proximity Duration | (R2) | Linear / Linear | ||||
| Image Grabs Variance | Spearman / Spearman | |||||
| Label Overrides Variance | Spearman / Pearson | |||||
@@ 4.4 Experience via Sensor-based Indicators
| Sensor Metric | Target Normality (Indiv. / Group) | Metric Normality (Indiv. / Group) | Correlation (Indiv. / Group) | p-Value (Indiv. / Group) | q-Value (FDR) (Indiv. / Group) | Test Used (Indiv. / Group) |
| Target: Presence (PQ + IPQ) | ||||||
| Shared Attention (%) | 0.8677 / 0.7494 | 0.3722 / 0.9363 | 0.0417 / 0.4802 | 0.7783 / 0.1141 | 0.8895 / 0.6085 | Pearson / Pearson |
| Jointly Viewed Images | 0.8677 / 0.7494 | 0.2048 / 0.3740 | -0.0820 / -0.4956 | 0.5795 / 0.1013 | 0.8895 / 0.6085 | Pearson / Pearson |
| Mean Pairwise Distance | 0.8677 / 0.7494 | 0.5384 / 0.7793 | 0.0736 / -0.0282 | 0.6190 / 0.9307 | 0.8895 / 0.9555 | Pearson / Pearson |
| Proximity Duration | 0.8677 / 0.7494 | 7.3267e-07 / 0.0650 | -0.0536 / 0.2626 | 0.7173 / 0.4097 | 0.8895 / 0.8582 | Spearman / Pearson |
| Variance in Image Grabs | 0.8677 / 0.7494 | 7.9738e-11 / 7.2687e-05 | 0.0083 / -0.1049 | 0.9555 / 0.7456 | 0.9555 / 0.8895 | Spearman / Spearman |
| Target: Cognitive Load (NASA TLX) | ||||||
| Shared Gaze Overlap | 0.0225 / 0.3828 | 7.3990e-08 / 0.0014 | 0.3223 / 0.2872 | 0.0255 / 0.3654 | 0.4080 / 0.8582 | Spearman / Spearman |
| Pairwise Distance | 0.0225 / 0.3828 | 0.5384 / 0.7793 | -0.1777 / -0.2020 | 0.2268 / 0.5289 | 0.8582 / 0.8895 | Spearman / Pearson |
| Variance in Image Grabs | 0.0225 / 0.3828 | 7.9738e-11 / 7.2687e-05 | 0.1261 / 0.2522 | 0.3933 / 0.4291 | 0.8582 / 0.8582 | Spearman / Spearman |
@@ 5 Discussion
Sensory behavioral features were weakly associated with subjective states such as the presence or cognitive load. A near-significant individual-level link was observed between cognitive load and shared gaze frequency, but it lacked strength and consistency at the group level. Presence did not correspond clearly with proximity or gaze patterns, suggesting a lack of visible manifestation of internal states in this context. Likewise, no significant relationship emerged between cognitive load and participation balance. H3 is partially accepted due to some individual-level signal, but not consistently. We reject H3.1, H3.2, and H3.3 due to a lack of strong evidence connecting presence or cognitive load to group-level behaviors. Our findings validate the reflective interaction mapping pathway, confirming it as a robust interpretive anchor in S2S framework. However, the lack of correlation in H3 identifies a perceptual blind spot where internal states like presence do not always leave a visible interaction trace in the Sensor Node. Although disappointing at first glance, this result replicates concerns by others [9, 25] that objective signals often fail to capture subjective experience unless paired with richer data sources.
@@ 5.1 Implications
Our results show that while sensed behavior reliably maps onto collaborative perception and task efficiency, it falls short of capturing deeper subjective experiences, such as presence or cognitive availability, in its current form. This reinforces that not all internal states are externally observable (3), even in high-fidelity sensing environments. This absence of mapping is not a flaw; it is a vital empirical finding that delineates what current MR sensing can (and importantly cannot) reveal about subjective experience. While prior works have treated gaze or physiological synchrony as proxies for presence or co-presence [28], our results underscore their limitations without deeper context. Our findings align with recent studies showing that even rich physiological or behavioral data often fail to reliably predict internal states unless complemented with richer modalities, such as neural, situational, or self-report inputs [41, 25, 29]. To meaningfully comprehend experience from sensor data, some intermediary techniques are needed to connect high-level behavioral indicators with low-level experiential states. Designers should treat behavioral signals as interpretive anchors, not definitive measures of internal states, reinforcing the cautious approach seen in immersive cognitive sensing research, and instead consider combining sensing with self-reports or adaptive prompting. Hybrid models that integrate system-logged data with lightweight subjective inputs could enable more robust personalization and responsiveness in collaborative MR.
@@ 6 Limitations and Future Work
This study offers foundational insight into the relationship between sensed group behavior, subjective perception, and task performance in collaborative MR settings. However, several limitations should be considered, some by design and scope, others presenting opportunities for future work.
First, we employed a singular task in one experimental condition to distinctly observe collaboration dynamics in a regulated environment. The open-ended image sorting task was chosen due to its flexibility and inherent coordination requirements, which are ideal for observing genuine group behavior. Despite this, the controlled environment may not completely reflect the full range of variability in time-sensitive or specialized tasks. This raises questions about generalizability, in terms of group size and across task types, spatial settings, and application domains. Furthermore, the technical requirement for single-user object ownership likely clustered group gaze toward active participants. While this mirrors a natural consensus-building workflow, it may lead to an overestimation of shared attention’s role compared to parallelized, multi-threaded tasks where users work independently. We view S2S as a foundational exploratory step; future confirmatory studies should address and assess various task types, multi-user object ownership, and adaptive system conditions to generalize beyond this setup.
@@
The number of groups restricts the statistical power of analyses conducted at the group level. Despite a relatively high participant count, the nature of small-group interactions limits group-level observations. Nonetheless, the sample size was adequate to reveal consistent and interpretable patterns across both individual and group analyses. Future research should build upon this by employing larger-scale deployments or multi-session studies. Group composition was not experimentally manipulated; participants had the autonomy to self-select or be randomly assigned to groups. This decision sought to observe genuine interaction dynamics instead of creating artificial teams. Nonetheless, pre-existing familiarity among participants might impact conversation, shared focus, and comfort. Although not measured in this study, future research should quantify prior familiarity to better understand its effects.
@@
We focused on interaction signals such as gaze, proximity, and voice activity, which can be easily tracked using standard MR hardware. This choice aligned with our aim to create practical, lightweight sensing systems, but using only headset sensors restricted us from accessing detailed affective, physiological, or facial data. Future research should investigate multimodal enhancements for richer sensing, such as emotion recognition or biometric monitoring. Finally, while our framework bridges sensor data and subjective reflection, it cannot fully capture internal states such as frustration, motivation, or attention lapses. These nuanced experiences often remain invisible to behavioral sensing alone. Our study highlights the gap between observable behavior and internal experience, reinforcing the need for interpretive unifying models that integrate multiple modalities and user perception.
In sum, the study was purposefully scoped to explore collaborative sensing in MR using a task and instrumentation that balance real-world validity and control. The observed patterns suggest promising directions for expanding system awareness and adaptability. Establishing a reliable alignment between sensor traces and subjective experience is a critical precursor to ”human-aware” MR systems. Practically, identifying these reflective anchors enables systems to transition from passive logging to active support, such as triggering automated ”collaboration nudges” (e.g., gaze-aware highlights to realign group focus) when the system detects a drop in shared attention. Theoretically, this mapping provides a verified lens to assess the quality of collaboration without requiring intrusive self-reports, allowing for more responsive and personalized collaborative environments . Future research should test these findings across various conditions, sensor modalities, and longer-term collaboration scenarios.
@@ 7 Conclusion
This study examines the alignment between objective sensor data and subjective perception of experiences in MR collaboration. The confirmation of H1 and H2 provides exploratory evidence that sensor-derived metrics reliably reflect perceived collaboration and task performance. The partial validation of H3 reveals limitations. Subjective states show limited correlation with objective measures of behavioral data, suggesting that internal experiences are difficult to capture solely through external observations. However, these results serve to inform the design of future confirmatory studies rather than provide a generalized predictive engine.
We situate this work in the growing body of research studying design-responsive, human-aware MR systems. By grounding system- and sensor-level observations in subjective experience, we move toward sensing models that prioritize users’ perspectives. This paper presents a conceptual model, a validated experimental setup, and empirical findings that illuminate where sensor data and user experience align and where they diverge. Through this lens, we take a step toward MR environments that support collaboration and sense it in the ways users feel it.
@@ Acknowledgements.
- [1] Apple Vision Pro. apple.com/apple-vision-pro/.
- [2] Hololens 2. learn.microsoft.com/en-us/hololens/.
@@
- [4] H. Bai, P. Sasikumar, J. Yang, and M. Billinghurst. A user study on mixed reality remote collaboration with eye gaze and hand gesture sharing. In Proceedings of CHI Conference on Human Factors in Computing Systems, 2020.
- [5] J. N. Bailenson, A. C. Beall, and J. Blascovich. Gaze and task performance in shared virtual environments. The Journal of Visualization and Computer Animation, 13(5), 2002.
- [6] J. N. Bailenson, A. C. Beall, J. Loomis, J. Blascovich, and M. Turk. Transformed social interaction: Decoupling representation from behavior and form in collaborative virtual environments. Presence: Teleoperators & Virtual Environments, 13(4), 2004.
- [7] J. N. Bailenson, N. Yee, J. Blascovich, A. C. Beall, N. Lundblad, and M. Jin. The use of immersive virtual reality in the learning sciences: Digital transformations of teachers, students, and social context. The Journal of the Learning Sciences, 17(1), 2008.
- [8] R. Bakeman and L. B. Adamson. Coordinating Attention to People and Objects in Mother-Infant and Peer-Infant Interaction. Wiley, Society for Research in Child Development, 55(4), 1984.
- [9] A. Bayro, Y. Ghasemi, and H. Jeong. Subjective and objective analyses of collaboration and co-presence in a virtual reality remote environment. In IEEE Conference on Virtual Reality and 3D User Interfaces Abstracts and Workshops, 2022.
@@
- [11] D. Bohus and E. Horvitz. Facilitating multiparty dialog with gaze, gesture, and speech. In ACM International Conference on Multimodal Interfaces and the Workshop on Machine Learning for Multimodal Interaction, 2010.
@@
- [39] M. R. Miller, H. Jun, and J. N. Bailenson. Motion and meaning: Sample-level nonlinear analyses of virtual reality tracking data. In IEEE International Symposium on Mixed and Augmented Reality Adjunct, 2021.
@@
- [41] N. Numan and A. Steed. Exploring User Behaviour in Asymmetric Collaborative Mixed Reality. In Proceedings of the 28th ACM Symposium on Virtual Reality Software and Technology, 2022.
@@
- [48] Z. Qu, R. Byrne, and M. Gorlatova. “Looking” into attention patterns in extended reality: An eye tracking-based study. In IEEE International Symposium on Mixed and Augmented Reality, 2024.
@@
- [50] J. A. Russell. A circumplex model of affect. Journal of Personality and Social Psychology, 39(6), 1980.
@@
- [67] G. Vissers and B. Dankbaar. Knowledge and Proximity. European Planning Studies, 21(5), 2013.
@@
- [69] J. Williamson, J. Li, V. Vinayagamoorthy, D. A. Shamma, and P. Cesar. Proxemics and social interactions in an instrumented virtual reality workshop. In Proceedings of the CHI Conference on Human Factors in Computing Systems, 2021.
@@
- [72] A. W. Woolley, C. F. Chabris, A. Pentland, N. Hashmi, and T. W. Malone. Evidence for a collective intelligence factor in the performance of human groups. Science, 330(6004), 2010.
- [73] Y. Yang, T. Dwyer, M. Wybrow, B. Lee, M. Cordeil, M. Billinghurst, and B. H. Thomas. Towards immersive collaborative sensemaking. Proceedings of the ACM on Human-Computer Interaction, 2022.
- [74] H. E. Yaremych and S. Persky. Tracing physical behavior in virtual reality: A narrative review of applications to social psychology. Journal of experimental social psychology, 85, 2019.